OpenStudy (ria23):

y=x^2-x-12 Convert to vertex form.

OpenStudy (ria23):

I got y-13=(x-1)^2 ;-; I don't think it's right...

OpenStudy (anonymous):

Well lets examine both Standard and Vertex form Standard \[f(x) = ax^2 + bx + c\] \[f(x) = a(x - h) + k\] (h,k) is the vertex so from standard form we can find the x-value of the vertex and then the y-value to find (h,k) and plug those into the first equation and a = 1 The vertex in standard form is \[x = \frac{ -b }{ 2a }\] Does this make sense so far?

OpenStudy (anonymous):

woops let me rewrite vertex form \[f(x) = a(x-h)^{2} + k\]

OpenStudy (ria23):

Yea, that makes sense.

OpenStudy (ria23):

I thought vertex form was written\[y-y _{1}=m(x-x _{1})^{2}\]

OpenStudy (anonymous):

Nah you are combining two different forms point-slop form for linear equations with vertex form

OpenStudy (anonymous):

so continuing on... our x-value of the vertex would be \[\frac{ -b }{ 2a } = \frac{ -(-1) }{ 2 }\] so our x-value is 1/2, now we got to plug that back into the original \[f(.5) = (.5)^2 - .5 - 12\] \[f(.5) = \frac{ 1 }{ 4 } - \frac{ 1 }{ 2 } - 12\] Get common denominators \[f(.5) = \frac{ 3 }{ 12 } - \frac{ 6 }{ 12 } - \frac{ 144 }{ 12 }\] \[f(.5) = \frac{ -147 }{ 12 }\] So our vertex is \[(\frac{ 1 }{ 2 }, \frac{ -147 }{ 12 })\] Now putting it into vertex form \[y = 1(x-\frac{ 1 }{ 2 })^{2} - \frac{ 147 }{ 12 }\]

OpenStudy (ria23):

Thank yhu for explaining everything. c:

OpenStudy (anonymous):

Yep!